# JWST/NIRSpec insights into the circumnuclear region of Arp 220: A detailed kinematic study

### v1

@@ 6.3 What drives the outflows?
The presence of AGN in the Arp 220 nuclei has been investigated in many previous works (e.g. Teng et al. 2015; Paggi et al. 2017; Yoast-Hull & Murray 2019; Perna et al. 2024), but conclusive evidence confirming their existence remains elusive. Furthermore, understanding whether outflows are powered by AGN or starburst presents an even greater challenge; in this section we relate the energetics of the Arp 220 multi-phase outflows to the properties of AGN (if present) and star formation, and we discuss possible scenarios.
@@
To evaluate the effect of the outflows on the star formation activity, we can compare their energetics with the star formation rate. However, the SFR varies depending on the band in which it is calculated, spanning values from 150 M⊙ yr-1 in the radio band (Varenius et al. 2016) to 0.5 M⊙ yr-1 through Pa (Perna et al. 2024) in the W nucleus, and from 80 to 0.17 M⊙ yr-1 in the E nucleus. Since the SFR measurements obtained from optical and near-infrared features may be significantly affected by dust attenuation, leading to underestimation, we focus on the case where the SFR is calculated using the radio band.
@@
Summing all the contributions of the outflows in the ionised and molecular phases in each nucleus, we derive a mass loading factor of 0.13 and 0.15 and a of erg s-1 and erg s-1 for the W and E nuclei, respectively. Note that these values are only marginally affected by the densities used, since the main contribution comes from the cold molecular gas. These values are similar to those measured in other local ULIRGs (Arribas et al., 2014; Chisholm et al., 2017; Lamperti et al., 2022) and high-z star-forming galaxies (Genzel et al. 2014; Heckman et al. 2015; Newman et al. 2012; Perna et al. 2017, 2018). Moreover, they are in agreement with models where feedback from supernovae is the main outflow driver (e.g. Finlator & Davé 2008; Heckman et al. 2015).
@@
Paggi et al. (2017) estimated AGN luminosities in the keV X-ray band of erg s-1 and erg s-1 for the W and E nuclei that lead to bolometric luminosities of erg s-1 and erg s-1, respectively. We also estimated the AGN luminosity from the total IR luminosity, assuming an AGN contribution of 6% to Lbol (Perna et al. 2021) and using the relative contribution of each nucleus measured in the ALMA continuum maps by Pereira-Santaella et al. (2021): L erg s-1 (W) and erg s-1 (E). It is worth noting that the AGN bolometric luminosity of the W nucleus inferred from FIR measurements is approximately 1 dex higher than the X-ray-based estimate, highlighting the challenges in constraining AGN activity in Arp 220. The outflow energetics computed in this section result in kinetic coupling efficiency values for the W nucleus, and for the W nucleus. These values are consistent with both theoretical predictions of kinetic coupling efficiencies in energy-conserving ( 0.005–0.05, e.g. Hopkins & Quataert 2010; Zubovas & King 2012; Costa et al. 2014; King & Pounds 2015) and radiation-pressure driven outflows (0.001–0.01, e.g. Costa et al. 2018; Ishibashi et al. 2018).
@@
These results show that there are no arguments at present to prefer SF- to AGN-driven outflows, since both of the scenario are possible. We note, however, that the typical dynamical timescales tdyn = Rout/ associated with the outflows in Arp 220 are approximately 1 Myr (see Table 3). In contrast, the SFR M⊙ yr-1 used to compare outflow energetics with expectations for starburst-driven outflows reflects the star formation history over a much longer period (up to 100 Myr; Kennicutt & Evans 2012). Following on these arguments, we consider two scenarios. In the former, we assume that the SFR in the last few million years is M⊙ yr-1. This would be consistent with the SFR measured in the host from UV, optical and near-infrared observations (e.g. Chandar et al. 2023; Perna et al. 2020, 2024; Giménez-Arteaga et al. 2022), as well as with the emission from supernovae and supernovae remnants at wavelengths from 2 to 18 cm which can be responsible for up to 20% of the total radio emission at GHz frequencies (but see the detailed discussion in Varenius et al. 2019). In this framework, the mass loading factors and outflow energetics would imply the presence of AGN winds, as SF would not provide enough energy to explain the outflow properties (, ). In the second scenario, we assume the presence of extremely high attenuation at the position of the two nuclei, of : such attenuation is required to match the SFR(PAH 3.3 m) 1 M⊙ yr-1 with measurements from FIR and GHz frequencies. With this assumption, both starburst- and AGN-driven winds can be responsible for the observed outflows, according to the outflow energetics computed in this work.
@@
In conclusion, further investigations are required to constrain the current SFR and the AGN activity in the nuclear regions of Arp 220. This information is essential for understanding the mechanisms driving the multi-phase and multi-scale outflows observed in this galaxy.
@@ 7 Conclusions
Using the exceptional JWST/NIRSpec facility, we investigated the complex multi-phase structure and kinematics in the nuclear region of the ULIRG merging system Arp 220. We spatially resolved the emission of ionized gas (Pa) and hot molecular gas (H2 1-0 S(1) 2.12 m) in the central kpc of the nuclear region of Arp 220, to identify and disentangle several high-velocity multi-phase outflows associated to the two nuclei of the system. We calculated the outflow mass and energetics of both ionized and hot molecular gas, and compared them with theoretical expectations for outflows driven by star formation and AGN activity. In the following, we summarize our main findings.
@@
We detected outflowing H2 and ionized gas with an hourglass-shape launched by the W nucleus, and extending out to a distance of 600 pc with velocities of 500 km s-1. We also found an ionized compact outflow around the W nucleus, with velocities up to 550 km s-1, without a hot molecular gas counterpart.
@@
We observed a hot molecular and ionized conical outflow in the south-west direction launched by the E nucleus, reaching velocities of up to 600 km s-1 and extending to the edges of the NIRSpec FoV (i.e. at least up to 400 pc). We also detected a multiphase blue-shifted, large scale ( 1 kpc) bubble in north-west direction with velocities up to 1100 km s-1. These high-velocity components, are aligned with the larger-scale ( kpc) outflow observed with MUSE, suggesting that they are part of the same outflow event originating from the E nucleus.
@@
We combined NIRSpec and ALMA outcomes to infer a hot-to-cold molecular gas ratio of for the outflows launched by the W and E nuclei, consistent with previous measurements for other ULIRGs from the literature. Considering together the material launched by the E and W nuclei, the largest contribution to the total outflow mass rate comes from cold molecular gas ( M⊙ yr-1), followed by ionized gas ( M⊙ yr-1, assuming [S ii]-based ne), and hot molecular gas ( 0.003 M⊙ yr-1).
@@
We discussed the possible origin of the outflows, and we concluded that there are no arguments at present to prefer SF- to AGN-driven outflow, since both of the scenario are possible. In particular, both recent nuclear activity with SFR M⊙ yr-1 and AGN luminosity of erg s-1 could explain the measured outflow mass loading factors (), mass outflow rates ( M⊙ yr-1) and kinetic powers ( erg s-1), according to theoretical predictions.
@@
NIRSpec observations allowed us to connect the 100 pc scale cold molecular outflows revealed by ALMA (e.g. Wheeler et al. 2020) with the kpc scale ionised and neutral outflows identified in the MUSE data cube (e.g. Perna et al. 2020). These findings highlight the significant role that individual outflows launched by each nucleus during a merger could play in the starburst-quasar evolutionary sequence: these outflows propagate in multiple directions, potentially impacting the overall physical and dynamical conditions of the host galaxy. Simulations predict that multi-directional outflows could lead to more uniform quenching of SF throughout the galaxy; in contrast, collimated outflows may be limited to specific directions and may allow SF to continue in other parts of the galaxy (see e.g. Wagner et al. 2012; Nelson et al. 2019; Torrey et al. 2020; Sivasankaran et al. 2024).
@@
We also investigated the properties of the ISM not participating in the outflow, in terms of recent SF traced by recombination lines, and obscuration. Additionally, we compared the stellar and the gas kinematics, both in the close surroundings of the two nuclei and across the large-scale disc. The key findings of this analysis are summarized below.
@@
We found, consistent with previous works, that SFR calculated with recombination lines ( 8 M⊙ yr-1) differs from the star formation computed with FIR and radio ( M⊙ yr-1). This result possibly indicates that a large fraction of the star formation is completely obscured at NIR wavelengths. Alternatively the FIR measurements of SFR may trace star-formation on different timescales.
@@
We estimated the attenuation from different emission line ratios: [Fe ii] 1.644 m/[Fe ii] 1.257 m, Br/Pa, and Pa/Pa. These values are consistent within 20%. Highest values are found at the position of the E nucleus, A10 mag; the lowest values (A mag) in the outer regions ( pc distance from the nuclei).
@@
Stellar and the gas kinematics show a relatively good match, with the stellar component following the ionized gas rotation and appearing to rotate faster than the hot molecular gas.
@@
The near-infrared, spatially resolved spectroscopic data of Arp 220 presented in this work demonstrates the power of JWST/NIRSpec to disentangle the complex kinematics in the nuclear regions, to understand the interplay between multi-phase outflows in the circumnuclear environments of nearby ULIRGs, and shed light on the feedback mechanisms in this peculiar phase of galaxy evolution.
Acknowledgements

### v2

@@ 6.2 The launching mechanism
The presence of AGN in the Arp 220 nuclei has been investigated in many previous works (e.g., Teng et al. 2015; Paggi et al. 2017; Yoast-Hull & Murray 2019; Perna et al. 2024), but conclusive evidence confirming their existence remains elusive. Furthermore, understanding whether outflows are powered by AGN or starburst presents an even greater challenge; in this section we relate the energetics of the Arp 220 multiphase outflows to the properties of AGN (if present) and SF, and we discuss possible scenarios.
@@
To evaluate the effect of the outflows on the SF activity, we can compare their energetics with the SFR. However, the SFR varies depending on the band in which it is calculated, spanning values from 150 M⊙ yr-1 in the radio band (Varenius et al. 2016) to 0.5 M⊙ yr-1 through Pa (Perna et al. 2024) in the W nucleus, and from 80 to 0.17 M⊙ yr-1 in the E nucleus. Since the SFR measurements obtained from optical and near-IR features may be significantly affected by dust attenuation, leading to underestimation, we focus on the case where the SFR is calculated using the radio band.
@@
Summing all the contributions of the outflows in the ionized and molecular phases in each nucleus, we derive a mass loading factor of 0.13 and 0.15 and a of erg s-1 and erg s-1 for the W and E nuclei, respectively. We note that these values are only marginally affected by the densities used, since the main contribution comes from the cold molecular gas. These values are similar to those measured in other local ULIRGs (Arribas et al., 2014; Chisholm et al., 2017; Lamperti et al., 2022) and high-z star-forming galaxies (Genzel et al. 2014; Heckman et al. 2015; Newman et al. 2012; Perna et al. 2017, 2018). Moreover, they are in agreement with models where feedback from supernovae is the main outflow driver (e.g., Finlator & Davé 2008; Heckman et al. 2015).
@@
Paggi et al. (2017) estimated AGN luminosities in the keV X-ray band of erg s-1 and erg s-1 for the W and E nuclei that lead to bolometric luminosities of erg s-1 and erg s-1, respectively. We also estimated the AGN luminosity from the total IR luminosity, assuming an AGN contribution of 6% to Lbol (Perna et al. 2021) and using the relative contribution of each nucleus measured in the ALMA continuum maps by Pereira-Santaella et al. (2021): L erg s-1 (W) and erg s-1 (E). It is worth noting that the AGN bolometric luminosity of the W nucleus inferred from FIR measurements is approximately 1 dex higher than the X-ray-based estimate, highlighting the challenges in constraining AGN activity in Arp 220. The outflow energetics computed in this section result in kinetic coupling efficiency values for the W nucleus, and for the W nucleus. These values are consistent with both theoretical predictions of kinetic coupling efficiencies in energy-conserving ( 0.005–0.05, e.g., Hopkins & Quataert 2010; Zubovas & King 2012; Costa et al. 2014; King & Pounds 2015) and radiation pressure-driven outflows (0.001–0.01, e.g., Costa et al. 2018; Ishibashi et al. 2018).
@@
These results show that there are no arguments at present to prefer SF- to AGN-driven outflows, since both of the scenario are possible. We note, however, that the typical dynamical timescales tdyn = Rout/ associated with the outflows in Arp 220 are approximately 1 Myr (see Table 3). In contrast, the SFR M⊙ yr-1 used to compare outflow energetics with expectations for starburst-driven outflows reflects the SF history over a much longer period (up to 100 Myr; Kennicutt & Evans 2012). Following on these arguments, we consider two scenarios. In the former, we assume that the SFR in the past few million years is M⊙ yr-1. This would be consistent with the SFR measured in the host from UV, optical and near-IR observations (e.g., Chandar et al. 2023; Perna et al. 2020, 2024; Giménez-Arteaga et al. 2022), as well as with the emission from supernovae and supernovae remnants at wavelengths from 2 to 18 cm which can be responsible for up to 20% of the total radio emission at GHz frequencies (but see the detailed discussion in Varenius et al. 2019). In this framework, the mass loading factors and outflow energetics would imply the presence of AGN winds, as SF would not provide enough energy to explain the outflow properties (, ). In the second scenario, we assume the presence of extremely high attenuation at the position of the two nuclei, of : such attenuation is required to match the SFR(PAH 3.3 m) 1 M⊙ yr-1 with measurements from FIR and GHz frequencies. With this assumption, both starburst- and AGN-driven winds can be responsible for the observed outflows, according to the outflow energetics computed in this work.
@@
In conclusion, further investigations are required to constrain the current SFR and the AGN activity in the nuclear regions of Arp 220. This information is essential for understanding the mechanisms driving the multiphase and multiscale outflows observed in this galaxy.
@@ 7 Conclusions
Using the exceptional JWST/NIRSpec facility, we investigated the complex multiphase structure and kinematics in the nuclear region of the ULIRG merging system Arp 220. We spatially resolved the emission of ionized gas (Pa) and hot molecular gas (H2 1-0 S(1) 2.12 m) in the central kiloparsec of the nuclear region of Arp 220 in order to identify and disentangle several high-velocity multiphase outflows associated with the two nuclei of the system. We modeled the emission of the outflows with and inferred the intrinsic geometric and kinematic properties, finding agreement with the properties inferred from standard methods. We calculated the outflow mass and energetics of both ionized and hot molecular gas, and compared them with theoretical expectations for outflows driven by SF and AGN activity. In the following, we summarize our main findings.
@@
We detected outflowing H2 and ionized gas with an hourglass shape launched by the W nucleus with an inclination of -60° with respect to the LOS as retrieved from and extending out to a distance of 600 pc with velocities of 500 km s-1. We also found an ionized compact outflow around the W nucleus, with velocities of up to 550 km s-1 and without a hot molecular gas counterpart.
@@
We observed a hot molecular and ionized conical outflow in the southwest direction launched by the E nucleus with an inclination of 70° with respect to the LOS reaching velocities of up to 600 km s-1 and extending to the edges of the NIRSpec FoV (i.e., at least up to 400 pc). We also detected a multiphase blueshifted large-scale ( 1 kpc) bubble in northwest direction with an inclination of 68° with respect to the LOS and with velocities of up to 1100 km s-1. These high-velocity components are aligned with the larger-scale ( kpc) outflow observed with MUSE, suggesting that they are part of the same outflow event originating from the E nucleus.
@@
We combined NIRSpec and ALMA outcomes to infer a hot-to-cold molecular gas ratio of for the outflows launched by the W and E nuclei, which is consistent with previous measurements for other ULIRGs from the literature. Considering the material launched by the E and W nuclei together, the largest contribution to the total outflow mass rate comes from cold molecular gas ( M⊙ yr-1), followed by ionized gas ( M⊙ yr-1, assuming [S ii]-based ne) and hot molecular gas ( 0.003 M⊙ yr-1).
@@
We discussed the possible origin of the outflows, and we conclude that there are no arguments at present to prefer SF- over AGN-driven outflows since both scenarios are possible. In particular, both recent nuclear activity with SFR M⊙ yr-1 and AGN luminosity of erg s-1 could explain the measured outflow mass loading factors (), mass outflow rates ( M⊙ yr-1), and kinetic powers ( erg s-1), according to theoretical predictions.
@@
The NIRSpec observations allowed us to connect the 100 pc scale cold molecular outflows revealed by ALMA (e.g., Wheeler et al. 2020) with the greater than kpc scale ionized and neutral outflows identified in the MUSE data cube (e.g., Perna et al. 2020). These findings highlight the significant role that individual outflows launched by each nucleus during a merger could play in the starburst-quasar evolutionary sequence. These outflows propagate in multiple directions on kiloparsec scales, potentially impacting the overall physical and dynamical conditions of the ISM in the host galaxy and are at odds with the more compact, spherical outflows detected in other nearby systems (e.g., Venturi et al. (2021)). Simulations predict that multidirectional outflows could lead to more uniform quenching of SF throughout the galaxy. In contrast, collimated outflows may be limited to specific directions and may allow SF to continue in other parts of the galaxy (see, e.g., Wagner et al. 2012; Nelson et al. 2019; Torrey et al. 2020; Sivasankaran et al. 2024).
@@
We also investigated the properties of the ISM not participating in the outflow in terms of recent SF traced by recombination lines and obscuration. Additionally, we compared the stellar and the gas kinematics in both the close surroundings of the two nuclei and across the large-scale disk. The key findings of this analysis are summarized below.
@@
Consistent with previous works, we found that SFR calculated with recombination lines ( 8 M⊙ yr-1) differs from the SF computed with FIR and radio ( M⊙ yr-1). This result possibly indicates that a large fraction of the SF is completely obscured at near-IR wavelengths. Alternatively, the FIR measurements of SFR may trace SF on different timescales.
@@
We estimated the attenuation from different emission line ratios: [Fe ii] 1.644 m/[Fe ii] 1.257 m, Br/Pa, and Pa/Pa. These values are consistent within 20%. The highest values are found at the position of the E nucleus, A10 mag, and the lowest values (A mag) are in the outer regions ( pc distance from the nuclei).
@@
The stellar and gas kinematics show a relatively good match, with the stellar component following the ionized gas rotation and appearing to rotate faster than the hot molecular gas.
@@
The near-IR, spatially resolved spectroscopic data of Arp 220 presented in this work demonstrates the power of JWST/NIRSpec to disentangle the complex kinematics in the nuclear regions and to understand the interplay between multiphase outflows in the circumnuclear environments of nearby ULIRGs. Moreover, the data shed light on the feedback mechanisms in this peculiar phase of galaxy evolution.
8 Data availability
The data cube, the spectra shown in Fig. 2, and line maps presented in Fig. 4 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/
Acknowledgements.